# Convergence for Natural Policy Gradient on Infinite-State Queueing MDPs

### v2

@@ 5.8. Rapid convergence for NPG on queueing MDPs
To see why, first recall from Section 3.3 that we require that the policy $\pi$ is non-idling: If the system state is not the all-zero state, the $\pi$ selects a service option which has a nonzero probability of completing a job. Recall also that on each time step, there’s a nonzero probability that no new job arrives. Thus, on each of the first $|z|$ time steps, if the system is not yet empty there is a nonzero probability that at least one job completes and no new job arrives. At some point during those steps, we will reach the all-zeros state.
@@ 6. Infinite State, Discounted Reward
In this section, we discuss the discounted setting: Rather than seeking to maximize the average reward, we seek to maximize the $\gamma$-discounted reward, for some $\gamma<1$, in expectation over some starting distribution $\eta$. In this setting, one can generalize existing finite-state results to the infinite-state setting. In particular, we can state the following theorem:
@@
In the infinite-state discounted-reward setting, if the discounted relative-value function $V^{\pi}_{\eta}$ is uniformly bounded over all policies $\pi$ for a given initial distribution $\eta$, then the NPG algorithm converges to the globally optimum discounted reward with $O(1/T)$ convergence rate.
@@
The proof of (Agarwal et al., 2021, Theorem 16) goes through nearly unchanged. The assumption of a finite state space is only to justify a bounded reward assumption, which in turn is only used in (Agarwal et al., 2021, Lemma 17), to prove that $V_{\eta}^{\pi}$, the expected discounted reward starting in initial-state-distribution $\eta$, is bounded. In the infinite-state setting, if $V^{\pi}_{\eta}$ is bounded, the proof still holds. ∎
This standard proof only uses its finite-state assumption to justify its bounded reward assumption. In particular, even when the state space is infinite, if the reward is bounded, the proof goes through unchanged. This is due in part to the fact that in the discounted-reward setting, a policy need not be stable to achieve finite reward.
In particular, if the reward is unbounded, but only grows polynomially from state to state, such as in the queueing setting, one can still prove that $V_{\eta}^{\pi}$ is bounded, and still prove a convergence-rate bound. More generally, it is sufficient for the worst-case discounted sum of rewards to be finite. The dependence on the discount rate $\gamma$ degrades in the $\gamma\to 1$ limit, depending on the growth rate of the reward, but the dependence on the number of iterations $T$ is unchanged.
Thus, in the discounted-reward setting, the standard proof of NPG’s convergence does generalize from the finite-state setting to the infinite-state setting, in contrast to the average-reward setting studied in this paper.
@@ 7. Conclusion
In the reinforcement learning (RL) setting, an important direction for future work would be to build off of our MDP optimization result to prove convergence-rate results for policy-gradient-based RL algorithms in the infinite-state setting.

### v3

@@ 5.8. Rapid convergence for NPG on queueing MDPs
To see why, first recall from Section 3.3 that our action space consists solely of non-idling actions: If the system state is not the all-zero state, the only service options in the action space are service options which have a nonzero probability of completing a job. Recall also that on each time step, there’s a nonzero probability that no new job arrives. Thus, on each of the first $|z|$ time steps, if the system is not yet empty there is a nonzero probability that at least one job completes and no new job arrives. At some point during those steps, we will reach the all-zeros state.
@@ 6. Infinite State, Discounted Reward
In this section, we discuss the discounted setting: Rather than seeking to maximize the average reward, we seek to maximize the $\gamma$-discounted reward, for some $\gamma<1$, in expectation over some starting distribution $\eta$. In this section, we demonstrate that existing finite-state results generalize to the infinite-state setting. In doing so, we remove Assumptions 1 and 2: They are not needed in the discounted setting, and we will specify the assumptions that are needed. Moreover, we remove any assumptions on the initial policy $\pi_{0}$: Any initialization suffices. We still work in the infinite-state-space, finite-action-space MDP setting specified in Section 3.1.
The $\gamma$-discounted reward $b^{\pi}_{\gamma}(\eta)$ under policy $\pi$ and initial state distribution $\eta$ is defined as follows:
(13) $b^{\pi}_{\gamma}(\eta)=\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}[r(s_{t},\pi)],$
where $s_{t}$ is a random variable denoting the state after $t$ steps under policy $\pi$ and initial state distribution $\eta$. The objective in the discounted setting is to find the policy $\pi$ which maximizes the discounted reward $b^{\pi}_{\gamma}(\eta)$ for a fixed initial state distribution $\eta$.
Note that $b^{\pi}_{\gamma}(\eta)$ may not converge for some policies $\pi$ or some initial distributions $\eta$. On the other hand, $b^{\pi}_{\gamma}(\eta)$ may converge even for unstable policies $\pi$: For instance, if $0\leq r(s,a)\leq 1$ for all states $s$ and actions $a$, then $0\leq b^{\pi}_{\gamma}(\eta)\leq\frac{1}{1-\gamma}$ for all policies $\pi$ and all initial state distributions $\eta$.
In this section, we start with the following assumption on $b^{\pi}_{\gamma}$:
Assumption 3.
We assume that for each initial state $s$, there exist upper and lower bounds $b_{\gamma}^{\min}(s)$ and $b_{\gamma}^{\max}(s)$ on the discounted reward of all policies $\pi$:
$\forall\pi,b_{\gamma}^{\min}(s)\leq b^{\pi}_{\gamma}(s)\leq b_{\gamma}^{\max}(s)$
We also define the discounted state visitation distribution $d^{\pi}_{\eta}$:
$d^{\pi}_{\eta}(s):=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s\mid\eta,\pi).$
Note that $d^{\pi}_{\eta}$ is well-defined for all policies $\pi$, including unstable policies.
We also make a second, slightly stronger assumption, namely that the expected discounted reward over any policy’s discounted state visitation distribution is also finite:
Assumption 4.
We assume that for each initial state $s$, there exist upper and lower bounds $c_{\gamma}^{\min}(s)$ and $c_{\gamma}^{\max}(s)$ on the expected discounted reward of any policy $\pi^{\prime}$ over the discounted state visitation distribution of any policy $\pi$, starting in state $s$:
$\forall\pi,\pi^{\prime},c_{\gamma}^{\min}(s)\leq\mathbb{E}[b^{\pi^{\prime}}_{\gamma}(d_{s}^{\pi})]\leq c_{\gamma}^{\max}(s)$
For example, both assumptions are satisfied by the GSSE MDPs defined in Section 3.3. Because at most $\ell n$ jobs can arrive per time step, $r(s_{t},\pi)$ is bounded by a linear function of $t$ for any policy $\pi$. This linear growth is overwhelmed by the exponential decay of $\gamma^{t}$, and thus (13) converges for any policy $\pi$, including unstable policies. This confirms Assumption 3. A similar argument also confirms Assumption 4 for GSSE MDPs, again using the linear bound on $r(s_{t},\pi)$ for all $\pi$.
With these assumptions, we demonstrate that (Agarwal et al., 2021, Theorem 16), a standard proof of convergence to optimality for NPG in the finite-state discounted-reward setting, generalizes to the infinite-state discounted-reward setting.
@@
(Generalization of (Agarwal et al., 2021, Theorem 16)) For any infinite-state finite-action discounted-reward MDP for which Assumptions 3 and 4 hold, the NPG algorithm achieves convergence rate $O(1/T)$ to the global optimum:
$b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{T}}_{\gamma}(\eta)\leq\frac{\ln A_{\max}}{\beta T}+\frac{\mathbb{E}[c^{\max}_{\gamma}(\eta)-c^{\min}_{\gamma}(\eta)]}{(1-\gamma)T}$
@@
We start by applying performance difference lemma to policies $\pi_{*}$, the optimal policy, and $\pi_{k}$, the $k$th NPG iterate. In the discounted reward setting, the performance difference lemma states:
(14) $b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{k}}_{\gamma}(\eta)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[Q_{\pi_{k}}(s,\pi^{*})-V_{\pi^{*}}(s)].$
Note that $b^{\pi_{*}}_{\gamma}(\eta)$ and $b^{\pi_{k}}_{\gamma}(\eta)$ are well-defined by Assumption 3, and note that the proof of the performance difference lemma (Agarwal et al., 2021, Lemma 2) holds in the infinite-state-space setting, even with unstable policies, because $d^{\pi_{*}}_{\eta}$ is well-defined whether or not $\pi_{*}$ is stable.
We now use the definition of the NPG algorithm in the discounted setting (Agarwal et al., 2021, Lemma 15):
$\pi_{k+1}(a\mid s)$ $=\pi_{k}(a\mid s)\beta^{(Q_{k}(s,a)-V_{k}(s))/(1-\gamma)}/Z_{s,k},$
where $\text{where }Z_{s,k}$ $=\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)\beta^{(Q_{k}(s,a^{\prime})-V_{k}(s))/(1-\gamma)}.$
Solving for $Q_{k}(s,a)$, we find that
$Q_{k}(s,a)-V_{k}(s)=\frac{1-\gamma}{\ln\beta}\ln\frac{\pi_{k+1}(a\mid s)Z_{s,k}}{\pi_{k}(a\mid s)}$
Substituting into (14), we have
$b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{k}}_{\gamma}(\eta)=\frac{1}{\ln\beta}\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[\sum_{a}\pi_{k+1}(a\mid s)\ln\frac{\pi_{k+1}(a\mid s)Z_{s,k}}{\pi_{k}(a\mid s)}]$
Rewriting in terms of the KL divergence, we have
(15) $b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{k}}_{\gamma}(\eta)=\frac{1}{\ln\beta}\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[KL(\pi_{*}(s|\cdot)\|\pi_{k}(s|\cdot))-KL(\pi_{*}(s|\cdot)\|\pi_{k+1}(s|\cdot))+\ln Z_{s,k}]$
Now, we apply (Agarwal et al., 2021, Lemma 17), which relates $Z_{s,k}$ to the change in discounted reward under the NPG policy. The proof of (Agarwal et al., 2021, Lemma 17) generalizes to the infinite-state setting without alteration. It states, using our notation, that for all distributions $\mu$:
(16) $b^{\pi_{k+1}}_{\gamma}(\mu)-b^{\pi_{k}}_{\gamma}(\mu)\geq\frac{1-\gamma}{\ln\beta}E_{s\sim\mu}[\ln Z_{s,k}]\geq 0$
Applying (16) and (15), and specifically the fact that $b^{\pi_{k+1}}_{\gamma}(\eta)\geq b^{\pi_{k}}_{\gamma}(\eta)$, by (16), we find:
$b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{T-1}}_{\gamma}(\eta)\leq\frac{1}{T}\sum_{t=0}^{T-1}(b^{\pi_{*}}_{\gamma}(\eta)-b^{\pi_{k}}_{\gamma}(\eta))$
$\leq\frac{1}{T\ln\beta}\sum_{t=0}^{T-1}\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[KL(\pi_{*}(s|\cdot)\|\pi_{k}(s|\cdot))-KL(\pi_{*}(s|\cdot)\|\pi_{k+1}(s|\cdot))+\ln Z_{s,k}]$
$\leq\frac{\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[KL(\pi_{*}(s|\cdot)\|\pi_{0}(s|\cdot))]}{T\ln\beta}+\frac{1}{(1-\gamma)T}\sum_{t=0}^{T-1}\mathbb{E}[b^{\pi_{k+1}}_{\gamma}(d^{\pi_{*}}_{\eta})-b^{\pi_{k}}_{\gamma}(d^{\pi_{*}}_{\eta})]$
$=\frac{\mathbb{E}_{s\sim d^{\pi_{*}}_{\eta}}[KL(\pi_{*}(s|\cdot)\|\pi_{0}(s|\cdot))]}{T\ln\beta}+\frac{\mathbb{E}[b^{\pi_{T}}_{\gamma}(d^{\pi_{*}}_{\eta})-b^{\pi_{0}}_{\gamma}(d^{\pi_{*}}_{\eta}])}{(1-\gamma)T}$
$\leq\frac{\ln A_{\max}}{T\ln\beta}+\frac{\mathbb{E}[b^{\pi_{T}}_{\gamma}(d^{\pi_{*}}_{\eta})-b^{\pi_{0}}_{\gamma}(d^{\pi_{*}}_{\eta})]}{(1-\gamma)T}$
$\leq\frac{\ln A_{\max}}{T\ln\beta}+\frac{\mathbb{E}[c^{\max}_{\gamma}(\eta)-c^{\min}_{\gamma}(\eta)]}{(1-\gamma)T}$
In the final step, we apply Assumption 4. Using the fact $b^{\pi_{T}}_{\gamma}(\eta)\geq b^{\pi_{T-1}}_{\gamma}(\eta)$ completes the proof. ∎
@@ 7. Conclusion
In the reinforcement learning (RL) setting, an important direction for future work would be to build off of our MDP optimization result to prove convergence-rate results for policy-gradient-based RL algorithms in the infinite-state setting. As important intermediate steps towards this goal, future work could study the NPG algorithm with function approximation in the infinite-state setting, as well as computationally-efficient variants of the NPG algorithm, such as those incorporating neural-network-based policy approximation.
8. Acknowledgments
We thank the area editor and the reviewers for their helpful comments and questions, which significantly improved the paper. We also acknowledge the following sources of funding: AFOSR Grant FA9550-24-1-0002, NSF Grants CNS 23-12714, CCF 22-07547, CNS 21-06801, EPCN-2144316, and CPS-2240982, the Tennenbaum Postdoctoral Fellowship at Georgia Tech, and a Northwestern University startup grant.